# Pseudo-Boolean Proof Logging for Optimal Classical Planning

### v1

@@ A PB circuit with input variables $V$ is a pair $\mathcal{R}=\langle R,r\rangle$, where
Abstraction heuristics such as pattern database (PDB) heuristics use goal distances in an induced abstract task for the heuristic estimates. Let $\Pi=\langle\mathcal{V},\mathcal{A},I,G\rangle$ be a STRIPS planning task. A PDB heuristic is defined in terms of a pattern $P\subseteq\mathcal{V}$ and its abstraction function $\alpha$ maps each state $s$ over $\mathcal{V}$ to an abstract state $\alpha(s)$ over $P$ as $\alpha(s)=s\cap P$. Each action $a\in\mathcal{A}$ induces the abstract action $a^{\alpha}$ with $\textit{pre}(a^{\alpha})=\textit{pre}(a)\cap P$, $\textit{add}(a^{\alpha})=\textit{add}(a)\cap P$, $\textit{del}(a^{\alpha})=\textit{del}(a)\cap P$, and $\textit{cost}(a^{\alpha})=\textit{cost}(a)$. The abstract goal $G^{\alpha}$ is $G\cap P$. The heuristic estimate of the PDB for state $s$ is the cost of an optimal solution of the abstract task $\langle P,\{a^{\alpha}\mid a\in\mathcal{A}\},\alpha(s),G^{\alpha}\rangle$ or $\infty$ if it is unsolvable.
Let $S_{\alpha}$ be the set of all abstract states. In practice, PDB heuristics do not build an abstract task for each heuristic evaluation but precompute the abstract goal distances $d(s_{\alpha})$ of all $s_{\alpha}\in S_{\alpha}$, and store them in a so-called pattern database. When a concrete state $s$ is evaluated, the PDB heuristic computes $\alpha(s)$ and returns the stored goal distance $d(\alpha(s))$ as the heuristic estimate $h(s)$.
For each abstract state $s_{\alpha}$, we will introduce a PB variable $r^{s_{\alpha}}$ that is true iff the variables from $\mathcal{V}$ encode a state $s$ with $\alpha(s)=s_{\alpha}$, and a variable $r^{s_{\alpha}}_{\geq B-d(s_{\alpha})}$ that is true iff the variables from $\mathcal{V}$ and $\mathcal{V}_{c}$ encode a pair $\langle s,c\rangle$ such that $\alpha(s)=s_{\alpha}$ and $c\geq\max\{B-d(s_{\alpha}),0\}$. The overall represented set contains all pairs $\langle s,c\rangle$ such that the abstract goal distance $d$ of $\alpha(s)$ is already so high that $c+d\geq B$, and thus it is impossible to reach the goal from $s$ with a strictly lower cost than $B$ if reaching $s$ already costs $c$.
Definition 5.
Let $\Pi=\langle\mathcal{V},\mathcal{A},I,G\rangle$ be a STRIPS planning task and $P\subseteq\mathcal{V}$. The PB circuit for the PDB heuristic with pattern $P$ is the PB circuit $\langle H_{\textup{PDB}},r_{\textup{PDB}}\rangle$ with input variables $\mathcal{V}\cup\mathcal{V}_{c}$, where $H_{\textup{PDB}}$ contains for each abstract state $s_{\alpha}\subseteq P$ the two constraints
$r^{s_{\alpha}}$ $\Leftrightarrow\sum\nolimits_{v\in s_{\alpha}}v+\sum\nolimits_{v\in P\setminus s_{\alpha}}\bar{v}\geq|P|$ (14)
and the final reification
@@
$(r_{s}\land\textit{cost}_{\geq\max\{0,B-h(s)\}})\rightarrow r_{\textup{PDB}}$ from $\mathcal{C}_{\textup{init}}\cup H_{\textup{PDB}}\cup\mathcal{C}_{s}$, where $\mathcal{C}_{s}=\{r_{s}\Leftrightarrow\sum_{v\in s}v+\sum_{v\in\mathcal{V}\setminus s}\bar{v}\geq|\mathcal{V}|\}$.
@@
$(r_{G}\land\overline{\textit{cost}_{\geq B}})\rightarrow\overline{r_{\textup{PDB}}}$ from $\mathcal{C}_{\textup{goal}}\cup H_{\textup{PDB}}$.
@@
The next lemma again considers individual actions and abstract states, but takes cost into account and derives that the successor situation is covered by the certificate.
@@
For each action $a\in\mathcal{A}$ and abstract state $s_{\alpha}$ such that $a^{\alpha}$ is applicable in $s_{\alpha}$, it is possible to derive
@@
We know that for the abstract goal distance it holds that $d(s_{\alpha}\llbracket a^{\alpha}\rrbracket)+\textit{cost}(a^{\alpha})\geq d(s_{\alpha})$. Together with $\textit{cost}(a^{\alpha})=\textit{cost}(a)$, we get $B-d(s_{\alpha}\llbracket a^{\alpha}\rrbracket)\leq B-d(s_{\alpha})+\textit{cost}(a)$. We use this to derive (b) $K^{\prime}_{\geq B-d(s_{\alpha})+\textit{cost}(a)}\rightarrow K^{\prime}_{\geq B-d(s_{\alpha}\llbracket a^{\alpha}\rrbracket)}$ as described in Lemma Proof-Logging Heuristic Search. In addition, we establish (c) $(r^{s_{\alpha}}\land r_{a})\rightarrow{r^{s_{\alpha}\llbracket a^{\alpha}\rrbracket}}^{\prime}$ by RUP (Lemma 5).
@@
The third lemma generalizes the previous lemma from individual actions to the entire transition relation. {restatable}lemmapdbinductivitylemmatwo For each abstract state $s_{\alpha}$ it is possible to derive $(r^{s_{\alpha}}_{\geq B-d(s_{\alpha})}\land r_{T})\rightarrow r^{\prime}_{\textup{PDB}}$ from $\mathcal{C}_{\textup{trans}}\cup H_{\textup{PDB}}\cup H_{\textup{PDB}}^{\prime}$.
@@
The final step for the inductivity lemma generalizes this from individual abstract states to the entire PB circuit of the heuristic. {restatable}lemmapdbinductivitylemma It is possible to derive the inductivity lemma $(r_{\textup{PDB}}\land r_{T})\rightarrow r^{\prime}_{\textup{PDB}}$ from $\mathcal{C}_{\textup{trans}}\cup H_{\textup{PDB}}\cup H_{\textup{PDB}}^{\prime}$.
@@
We considered PDB heuristics as one example of abstractions because they have a particularly simply structured abstraction function. To adapt the approach to other abstraction heuristics, the overall line of argument still works but we would have to replace (14) with one or more reifications that allow to identify the concrete states that correspond to a given abstract state. In addition, the derivations from Lemmas 5, 5 and 5 need to be adapted accordingly.
The described proof logging only has a constant-factor overhead for a somewhat naive implementation of PDBs. While the state and goal lemma only require a single RUP statement, the inductivity lemma requires effort in $O(|S_{\alpha}||\mathcal{A}|)$, iterating once over all actions for each abstract state. A better PDB implementation can deal more efficiently with abstract states that are not reverse-reachable. In the supplementary material we describe a variant of the certificate that does not explicitly represent such states and which guarantees a constant-factor overhead also for such a more efficient computation of PDBs.
@@
The maximum heuristic (Bonet and Geffner 2001) $h^{\textup{max}}(s)$ is defined as $h^{\textup{max}}(s,G)$, where $h^{\textup{max}}(s,V)$ is the pointwise greatest function with $h^{\textup{max}}(s,V)=\begin{cases}0&\text{if $V\subseteq s$}\\ \min\limits_{\begin{subarray}{c}a\in\mathcal{A},\\ v\in\textit{add}(a)\end{subarray}}(\textit{cost}(a)+h^{\textup{max}}(s,\textit{pre}(a)))&\text{if $|V|=1$ and $V\not\subseteq s$}\\ \max\limits_{v\in V}h^{\textup{max}}(s,\{v\})&\text{otherwise}\end{cases}$
For $v\in\mathcal{V}$, the max value is $V^{\textup{max}}(v)=h^{\textup{max}}(s,\{v\})$. A typical efficient implementation of the maximum heuristic determines this value for all $v$ with $V^{\textup{max}}(v)<h^{\textup{max}}(s)$. We define $W^{\textup{max}}(v)=\min\{V^{\textup{max}}(v),h^{\textup{max}}(s)\}$.
@@
$r_{v}\Leftrightarrow\bar{v}+K_{\geq B-h^{\textup{max}}(s)+W^{\textup{max}}(v)}\geq 1$ (17)
$r^{\textup{max}}_{s}\Leftrightarrow\sum_{v\in\mathcal{V}}r_{v}\geq|\mathcal{V}|$ (18)
The PB circuit for state $s$ then is $\langle H_{\textup{max}},r^{\textup{max}}_{s}\rangle$, where $H_{\textup{max}}$ is the sequence of all reifications (17) and (18).
@@
$(r_{s}\land\textit{cost}_{\geq\max\{0,B-h^{\textup{max}}(s)\}})\rightarrow r^{\textup{max}}_{s}$ from $\mathcal{C}_{\textup{init}}\cup H_{\textup{max}}\cup\mathcal{C}_{s}$, where $\mathcal{C}_{s}=\{r_{s}\Leftrightarrow\sum_{v\in s}v+\sum_{v\in\mathcal{V}\setminus s}\bar{v}\geq|\mathcal{V}|\}$.
@@
$(r_{G}\land\overline{\textit{cost}_{\geq B}})\rightarrow\overline{r^{\textup{max}}_{s}}$ from $\mathcal{C}_{\textup{goal}}\cup H_{\textup{max}}$.
@@
$(r^{\textup{max}}_{s}\land r_{a})\rightarrow{r^{\textup{max}}_{s}}^{\prime}$ from $\mathcal{C}_{\textup{trans}}\cup H_{\textup{max}}\cup H_{\textup{max}}^{\prime}$.
@@
If $W^{\textup{max}}(p)=h^{\textup{max}}(s)$, we exploit that the $h^{\textup{max}}$ cost of $\textit{pre}(a)\geq B$ is so high that $K_{\geq B}\geq 1$, which transfers to the primed cost.
If $W^{\textup{max}}(p)<h^{\textup{max}}(s)$ we derive $(r^{\textup{max}}_{s}\land r_{a})\rightarrow r^{\prime}_{v}$ for all variables $v$ by RUP. If $v\in\textit{del}(a)$, ($\overline{v^{\prime}}\geq 1$), then $r^{\prime}_{v}\geq 1$, independent of the cost. If $v\in\textit{add}(a)$, we exploit that $W^{\textup{max}}(p)=V^{\textup{max}}(p)$ and consequently $h^{\textup{max}}(s,\textit{pre}(a))=V^{\textup{max}}(p)$ because $p$ has maximum $V^{\textup{max}}$-value among $a$’s preconditions. This implies that that $V^{\textup{max}}(v)\leq V^{\textup{max}}(p)+\textit{cost}(a)$ and consequently $W^{\textup{max}}(v)\leq W^{\textup{max}}(p)+\textit{cost}(a)$. We can derive that $K^{\prime}_{\geq B-h^{\textup{max}}(s)+W^{\textup{max}}(v)}\geq 1$, which is sufficient to derive $r_{v}\geq 1$ also in this case. If $v\notin\textit{evars}(a)$, we can derive from assumption $\overline{r^{\prime}_{v}}\geq 1$ that $v^{\prime}\geq 1$ and $\overline{K^{\prime}_{\geq B-h^{\textup{max}}(s)+W^{\textup{max}}(v)}}\geq 1$. Both properties can be propagated to the unprimed variables within some steps, contradicting the assumption $r^{\textup{max}}_{s}\geq 1$. With the subresults $(r^{\textup{max}}_{s}\land r_{a})\rightarrow r^{\prime}_{v}$, RUP can derive that $(r^{\textup{max}}_{s}\land r_{a})\rightarrow{r^{\textup{max}}_{s}}^{\prime}$.
The proof for an empty precondition is analogous without the need to use $p$ to establish a certain $K_{\geq\cdot}$. ∎
@@
$(r^{\textup{max}}_{s}\land r_{T})\rightarrow{r^{\textup{max}}_{s}}^{\prime}$ from $\mathcal{C}_{\textup{trans}}\cup H_{\textup{max}}\cup H_{\textup{max}}^{\prime}$.
@@
Extending $h^{\textup{max}}$ with proof logging only leads to the following overhead: adding reifications (17) for each $v\in\mathcal{V}$ and (18) takes additional time linear in $|\mathcal{V}|$ because $h(s)$ and the values necessary for $W^{\textup{max}}{}$ are already computed during the normal evaluation. In addition, it has to reinitialize some values for each variable before each new evaluation, so the overhead is within a constant factor.

### v2

@@ A PB circuit with input variables $V$ is a pair $\mathcal{R}=\langle R,r\rangle$, where
Abstraction heuristics such as pattern database (PDB) heuristics use goal distances in an induced abstract task for the heuristic estimates. Let $\Pi=\langle\mathcal{V},\mathcal{A},I,G\rangle$ be a STRIPS planning task. A PDB heuristic is defined in terms of a pattern $P\subseteq\mathcal{V}$ and its abstraction function $\alpha$ maps each state $s$ over $\mathcal{V}$ to an abstract state $\alpha(s)$ over $P$ as $\alpha(s)=s\cap P$. Each action $a\in\mathcal{A}$ induces the abstract action $a^{\alpha}$ with $\textit{pre}(a^{\alpha})=\textit{pre}(a)\cap P$, $\textit{add}(a^{\alpha})=\textit{add}(a)\cap P$, $\textit{del}(a^{\alpha})=\textit{del}(a)\cap P$, and $\textit{cost}(a^{\alpha})=\textit{cost}(a)$. The abstract goal $G^{\alpha}$ is $G\cap P$. The heuristic estimate of the PDB for state $s$ is the cost of an optimal solution of the abstract task $\langle P,\{a^{\alpha}\mid a\in\mathcal{A}\},\alpha(s),G^{\alpha}\rangle$ or $\infty$ if it is unsolvable. This is a lower bound of the goal distance of $s$ in $\Pi$, as any solution of the concrete task corresponds to a solution of the abstract task with equal cost. In practice, a PDB heuristic does not build an abstract task for each heuristic evaluation but precomputes the abstract goal distances $d(s_{\alpha})$ for all abstract states $s_{\alpha}$ and stores them in a so-called pattern database. When a concrete state $s$ is evaluated, the PDB heuristic computes $\alpha(s)$ and returns the stored goal distance $d(\alpha(s))$ as the heuristic estimate $h(s)$.
The invariant for a PDB heuristic should hold for all pairs $\langle s,c\rangle$ such that the abstract goal distance $d$ of $\alpha(s)$ is already so high that $c+d\geq B$, and thus it is impossible to reach the goal from $s$ with a strictly lower cost than $B$ if reaching $s$ already costs $c$.
Let $S_{\alpha}$ be the set of all abstract states. For each $s_{\alpha}\in S_{\alpha}$, the heuristic introduces two PB variables. The variable $r^{s_{\alpha}}$ is true iff the variables from $\mathcal{V}$ encode a state $s$ with $\alpha(s)=s_{\alpha}$ by adding a reification
$r^{s_{\alpha}}\Leftrightarrow\sum\nolimits_{v\in s_{\alpha}}v+\sum\nolimits_{v\in P\setminus s_{\alpha}}\bar{v}\geq|P|.$ (14)
The variable $r^{s_{\alpha}}_{\geq B-d(s_{\alpha})}$ is true iff the variables from $\mathcal{V}$ and $\mathcal{V}_{c}$ encode a pair $\langle s,c\rangle$ such that $\alpha(s)=s_{\alpha}$ and $c\geq\max\{B-d(s_{\alpha}),0\}$ by adding reification
$r^{s_{\alpha}}_{\geq B-d(s_{\alpha})}\Leftrightarrow r^{s_{\alpha}}+K_{\geq B-d({s_{\alpha}})}\geq 2.$ (15)
As a final reification for the invariant the heuristic adds
@@
The PB circuit for any state then is $\langle H_{\textup{PDB}},r_{\textup{PDB}}\rangle$, where $H_{\textup{PDB}}$ is the sequence of all reifications (14), (15) and (16). On the evaluation of a state $s$, the heuristic always returns reification variable $r_{\textup{PDB}}$ to the search. In the following we discuss how the required proofs can be generated. This generation relies on the correctness and admissibility of the PDB heuristic, but the generated proofs themselves do not.
@@
$(r_{s}\land\textit{cost}_{\geq\max\{0,B-h(s)\}})\rightarrow r_{\textup{PDB}}$ from $\mathcal{C}_{s}\cup H_{\textup{PDB}}\cup\mathcal{C}_{\geq}\cup\mathcal{C}_{K}$, where $\mathcal{C}_{s}=\{r_{s}\Leftrightarrow\sum_{v\in s}v+\sum_{v\in\mathcal{V}\setminus s}\bar{v}\geq|\mathcal{V}|\}$.
@@
$(r_{G}\land{r_{\textup{PDB}}})\rightarrow{\textit{cost}_{\geq B}}$ from $\mathcal{C}_{\textup{goal}}\cup H_{\textup{PDB}}\cup\mathcal{C}_{\geq}\cup\mathcal{C}_{K}$.
@@
$(r^{s_{\alpha}}\land r_{a})\rightarrow{r^{s_{\alpha}\llbracket a^{\alpha}\rrbracket}}^{\prime}$ from $\mathcal{C}_{\textup{trans}}\cup H_{\textup{PDB}}\cup H_{\textup{PDB}}^{\prime}\cup\mathcal{C}_{\geq}\cup\mathcal{C}_{K}.$
The next lemma again considers individual actions and abstract states, but takes cost into account and derives that the invariant of the certificate is true for the successor state-cost pair.
@@
For each action $a\in\mathcal{A}$ and abstract state $s_{\alpha}$ such that $a^{\alpha}$ is applicable in $s_{\alpha}$, it is possible to derive $(r^{s_{\alpha}}_{\geq B-d(s_{\alpha})}\land r_{a})\rightarrow r^{\prime}_{\textup{PDB}}$ from $\mathcal{C}_{\textup{trans}}\cup H_{\textup{PDB}}\cup H_{\textup{PDB}}^{\prime}\cup\mathcal{C}_{\geq}\cup\mathcal{C}_{K}$.
@@
We know that for the abstract goal distance it holds that $d(s_{\alpha}\llbracket a^{\alpha}\rrbracket)+\textit{cost}(a^{\alpha})\geq d(s_{\alpha})$. Together with $\textit{cost}(a^{\alpha})=\textit{cost}(a)$, we get $B-d(s_{\alpha}\llbracket a^{\alpha}\rrbracket)\leq B-d(s_{\alpha})+\textit{cost}(a)$. We use this to derive (b) $K^{\prime}_{\geq B-d(s_{\alpha})+\textit{cost}(a)}\rightarrow K^{\prime}_{\geq B-d(s_{\alpha}\llbracket a^{\alpha}\rrbracket)}$ as described in Lemma Proof-Logging Heuristic Search. In addition, we establish (c) $(r^{s_{\alpha}}\land r_{a})\rightarrow{r^{s_{\alpha}\llbracket a^{\alpha}\rrbracket}}^{\prime}$ by RUP (Lemma Proof-Logging Pattern Database Heuristics).
@@
The third lemma generalizes the previous lemma from individual actions to the entire transition relation. {restatable}lemmapdbinductivitylemmatwo For each abstract state $s_{\alpha}$ it is possible to derive $(r^{s_{\alpha}}_{\geq B-d(s_{\alpha})}\land r_{T})\rightarrow r^{\prime}_{\textup{PDB}}$ from $\mathcal{C}_{\textup{trans}}\cup H_{\textup{PDB}}\cup H_{\textup{PDB}}^{\prime}\cup\mathcal{C}_{\geq}\cup\mathcal{C}_{K}$.
@@
The final step for the inductivity lemma generalizes this from individual abstract states to the entire PB circuit of the heuristic. {restatable}lemmapdbinductivitylemma It is possible to derive the inductivity lemma $(r_{\textup{PDB}}\land r_{T})\rightarrow r^{\prime}_{\textup{PDB}}$ from $\mathcal{C}_{\textup{trans}}\cup H_{\textup{PDB}}\cup H_{\textup{PDB}}^{\prime}\cup\mathcal{C}_{\geq}\cup\mathcal{C}_{K}$.
@@
We considered PDB heuristics as one example of abstractions because they have a particularly simply structured abstraction function. To adapt the approach to other abstraction heuristics, the overall line of argument still works but we would have to replace (14) with one or more reifications that allow to identify the concrete states that correspond to a given abstract state. In addition, the derivations from Lemmas Proof-Logging Pattern Database Heuristics, Proof-Logging Pattern Database Heuristics and Proof-Logging Pattern Database Heuristics need to be adapted accordingly.
The described proof logging only has a constant-factor overhead for a somewhat naive implementation of PDBs. While the state and goal lemma only require a single RUP statement, the inductivity lemma requires effort in $O(|S_{\alpha}||\mathcal{A}|)$, iterating once over all actions for each abstract state. A better PDB implementation can deal more efficiently with abstract states that are not reverse-reachable. In the appendix we describe a variant of the certificate that does not explicitly represent such states and which guarantees a constant-factor overhead also for such a more efficient computation of PDBs.
@@
[…]
For $v\in\mathcal{V}$, the max value is $V^{\textup{max}}(v)=h^{\textup{max}}(s,\{v\})$, which can be seen as the cost of making $v$ true (starting from $s$) under this relaxation. The heuristic certificate will build on the following insight: since $h^{\textup{max}}$ is admissible, it will be impossible to reach the goal from $s$ with overall cost $<B$ if the cost to reach $s$ already exceeds $B-h^{\textup{max}}(s)$. If in a state $\hat{s}$ additional variables are true, the state can be closer to the goal but under the relaxation of the heuristic the advantage gained from an individual variable $v$ will be at most $V^{\textup{max}}(v)$, so for arbitrary non-empty states $\hat{s}$, the goal cannot be reached if the cost to reach $\hat{s}$ already exceeds $B-h^{\textup{max}}(s)+\max_{v\in\hat{s}}V^{\textup{max}}(v)$ (note that $V^{\textup{max}}(v)=0$ for $v\in s$), even under the relaxation of the heuristic. In the delete-relaxed task, the empty state cannot be closer to the goal than $s$, so we can use bound $B-h^{\textup{max}}(s)$.
A typical efficient implementation of the maximum heuristic only determines the max value for all $v$ with $V^{\textup{max}}(v)<h^{\textup{max}}(s)$. We can still build a valid heuristic certificate, requiring that the cost $B$ has already been exceeded for states that contain a variable for which $V^{\textup{max}}(v)\geq h^{\textup{max}}(s)$. For the formal definition of the certificate, we will handle this aspect by an auxiliary value $W^{\textup{max}}(v)=\min\{V^{\textup{max}}(v),h^{\textup{max}}(s)\}$.
@@
$r_{v,s}\Leftrightarrow\bar{v}+K_{\geq B-h^{\textup{max}}(s)+W^{\textup{max}}(v)}\geq 1$ (17)
In addition, it adds the following reification:
$r^{\textup{max}}_{s}\Leftrightarrow K_{\geq B-h^{\textup{max}}(s)}+\sum_{v\in\mathcal{V}}r_{v,s}\geq|\mathcal{V}|+1$ (18)
The PB circuit for state $s$ then is $\langle H_{\textup{max},s},r^{\textup{max}}_{s}\rangle$, where $H_{\textup{max},s}$ is the sequence of all reifications (17) and (18).
Intuitively, (17) expresses the implication that the cost is at least $B-h^{\textup{max}}(s)+W^{\textup{max}}(v)$ if $v$ is true. Equation (18) corresponds to a conjunction of $K_{\geq B-h^{\textup{max}}(s)}$ and such implications for each variable, so overall $r^{\textup{max}}_{s}$ will be true for all state value pairs $\langle\hat{s},\hat{c}\rangle$, where $\hat{s}\neq\emptyset$ and $\hat{c}\geq\max_{v\in\hat{s}}\{B-h^{\textup{max}}(s)+W^{\textup{max}}(v)\}$ or where $\hat{s}=\emptyset$ and $\hat{c}\geq B-h^{\textup{max}}(s)$.
@@
$(r_{s}\land\textit{cost}_{\geq\max\{0,B-h^{\textup{max}}(s)\}})\rightarrow r^{\textup{max}}_{s}$ from $\mathcal{C}_{s}\cup H_{\textup{max},s}\cup\mathcal{C}_{\geq}\cup\mathcal{C}_{K}$, where $\mathcal{C}_{s}=\{r_{s}\Leftrightarrow\sum_{v\in s}v+\sum_{v\in\mathcal{V}\setminus s}\bar{v}\geq|\mathcal{V}|\}$.
@@
$(r_{G}\land{r^{\textup{max}}_{s}})\rightarrow{\textit{cost}_{\geq B}}$ from $\mathcal{C}_{\textup{goal}}\cup H_{\textup{max},s}\cup\mathcal{C}_{\geq}\cup\mathcal{C}_{K}$.
@@
$(r^{\textup{max}}_{s}\land r_{a})\rightarrow{r^{\textup{max}}_{s}}^{\prime}$ from $\mathcal{C}_{\textup{trans}}\cup H_{\textup{max},s}\cup H_{\textup{max},s}^{\prime}\cup\mathcal{C}_{\geq}\cup\mathcal{C}_{K}$.
@@
The overall argument of the proof is as follows:
If the action has a precondition $p$, we can conclude that $p\geq 1$, which allows us to derive with (17) that $K_{\geq B-h^{\textup{max}}(s)+W^{\textup{max}}(p)}\geq 1$. We can establish $K^{\prime}_{\geq B-h^{\textup{max}}(s)}$ using Lemmas Proof-Logging Heuristic Search and Proof-Logging Heuristic Search.
To establish $r^{\prime}_{v,s}\geq 1$ for all variables, we consider the precondition with maximal $W^{\textup{max}}$ value.
If $W^{\textup{max}}(p)=h^{\textup{max}}(s)$, then $K_{\geq B-h^{\textup{max}}(s)+W^{\textup{max}}(p)}\geq 1$ corresponds to $K_{\geq B}\geq 1$. We derive $K^{\prime}_{\geq B}\geq 1$, which is sufficient to establish $r^{\prime}_{v,s}\geq 1$.
If $W^{\textup{max}}(p)<h^{\textup{max}}(s)$, we establish $r^{\prime}_{v,s}\geq 1$ depending on whether and how $v$ occurs in the effect of $a$. If it is a delete effect, then $\overline{v^{\prime}}\geq 1$. If it is an add effect, we exploit that $p$ has maximal $W^{\textup{max}}$ value among the preconditions, so from the definition of the maximum heuristic we can see that $W^{\textup{max}}(v)\leq W^{\textup{max}}(p)+\textit{cost}(a)$. Establishing $K^{\prime}_{\geq B-h^{\textup{max}}(s)+W^{\textup{max}}(p)+\textit{cost}(a)}$ and using this insight gives $K^{\prime}_{\geq B-h^{\textup{max}}(s)+W^{\textup{max}}(v)}\geq 1$. If $v\notin\textit{evars}(a)$, we can carry over the reason for $r_{v,s}\geq 1$ to the primed variables.
If the precondition of the action is empty, we build on $K_{\geq B-h^{\textup{max}}(s)}$, which gives $K^{\prime}_{\geq B-h^{\textup{max}}(s)+\textit{cost}(a)}$ and $K^{\prime}_{\geq B-h^{\textup{max}}(s)}$ using Lemmas Proof-Logging Heuristic Search and Proof-Logging Heuristic Search. For establishing $r^{\prime}_{v,s}\geq 1$ for all variables, we proceed as before for the case $W^{\textup{max}}(p)<h^{\textup{max}}(s)$ (replacing $W^{\textup{max}}(p)$ with 00). For the special case $h^{\textup{max}}(s)=0$, we can instead use the shorter argument as in the previous case $W^{\textup{max}}(p)=h^{\textup{max}}(s)$, building on $K_{\geq B}\geq 1$.
The technical details are included in the appendix. ∎
@@
$(r^{\textup{max}}_{s}\land r_{T})\rightarrow{r^{\textup{max}}_{s}}^{\prime}$ from $\mathcal{C}_{\textup{trans}}\cup H_{\textup{max},s}\cup H_{\textup{max},s}^{\prime}\cup\mathcal{C}_{\geq}\cup\mathcal{C}_{K}$.
@@
Extending $h^{\textup{max}}$ with proof logging only leads to the following overhead: adding reifications (17) for each $v\in\mathcal{V}$ and (18) takes additional time linear in $|\mathcal{V}|$ because $h(s)$ and the values necessary for $W^{\textup{max}}{}$ are already computed during the normal evaluation. In addition, the implementation has to reinitialize some values for each variable before each new evaluation, so the overhead is within a constant factor.